# Accelerated Decay due to Operator Spreading in Bulk-Dissipated Quantum Systems

### v2

@@ Abstract
which is expressed as ~*$\tilde{\mathcal{L}}^{*}\hat{A}=\mathcal{L}(\hat{A}\rho_{\mathrm{ss}})\rho_{\mathrm{ss}}^{-1}$ [17]. We assume that $\rho_{\mathrm{ss}}$ is invertible for simplicity. The symmetrized Liouvillian is a non-positive and Hermitian superoperator (i.e., ~$\tilde{\mathcal{L}}_{s}\leq 0$ and ~s*=~$\tilde{\mathcal{L}}_{s}^{*}=\tilde{\mathcal{L}}_{s}$). The non-positivity is shown by using the complete positivity of $\mathcal{L}$. We sort the eigenvalues of ~$\tilde{\mathcal{L}}_{s}$ denoted by $\{s_{\alpha}\}$ in the descending order: $0=s_{0}\geq s_{1}\geq s_{2}\geq\dots$. ~$\tilde{\mathcal{L}}_{s}$ has a zero eigenvalue: ~$\tilde{\mathcal{L}}_{s}\hat{I}=0$, where $\hat{I}$ is the identity operator. The spectral gap $g_{s}$ of ~$\tilde{\mathcal{L}}_{s}$ is defined by
@@
We assume weak dissipation and set $\gamma=0.01$. This open quantum system has been implemented on Rydberg atoms under laser driving and dissipation [20, 21]. Each Rydberg atom is regarded as a two-level system, where the spin up (down) corresponds to the Rydberg (ground) state. In this model, the Liouvillian gap is independent of the system size (i.e., $g\sim O(L^{0})$. See Fig. 1.).
@@
$\kappa_{A}^{(0)}(t)=\frac{\braket{\hat{A}_{0}(t),\tilde{\mathcal{L}}_{s}\hat{A}_{0}(t)}_{\mathrm{ss}}}{\braket{\hat{A}_{0}(t),\hat{A}_{0}(t)}_{\mathrm{ss}}},\quad\hat{A}_{0}(t)=e^{i\hat{H}t}\hat{A}e^{-i\hat{H}t}.$ (20)
@@
$\frac{d\kappa_{A}(t)}{dt}=\dot{\kappa}_{A}-2\delta\kappa_{A}^{2},$ (21)
@@
$\dot{\kappa}_{A}$ $=\frac{\braket{\hat{A}(t),[\tilde{\mathcal{L}},\tilde{\mathcal{L}}_{s}]\hat{A}(t)}_{\mathrm{ss}}}{\braket{\hat{A}(t),\hat{A}(t)}_{\mathrm{ss}}},$
@@
This formula generalizes the result in [25] to any open quantum systems obeying Lindblad equations. $\dot{\kappa}_{A}/\gamma$ represents how fast the operator size of $A(t)$ spreads with time and $(\delta\kappa_{A}/\gamma)^{2}$ roughly gives the variance of the operator size. Note that the time evolution of $A(t)$ in $\dot{\kappa}_{A}$ is well approximated by the unitary dynamics without dissipation [18].
The formula accurately explains the observed three dynamic regimes in Fig. 1. In the acceleration regime, recent studies on random unitary circuits [23, 24] indicate $\dot{\kappa}_{A}\sim\gamma$ and $\delta\kappa_{A}^{2}\sim\gamma^{2}t$. Since $\dot{\kappa}_{A}\gg\delta\kappa_{A}^{2}$ for small $\gamma$,
@@
The system enters the plateau regime when the operator spreads across the entire system at $t\sim N$. Accordingly, $\kappa_{A}(t)$ is proportional to $\gamma N$. In this regime, $\dot{\kappa}_{A}\approx 0$, whereas $\delta\kappa_{A}^{2}\sim\gamma^{2}N$ [18]. Thus, during the duration time $\Delta t$ in the plateau regime, $\kappa_{A}(t)$ reduces by the amount proportional to $\gamma^{2}N\Delta t$. This decrease is negligible compared with the plateau value $\sim\gamma N$ up to $\Delta t\sim\gamma^{-1}$. We therefore conclude that $\kappa_{A}(t)$ approximately remains constant up to $t\sim\gamma^{-1}$:
@@
Here, we briefly explain the dynamics of $\kappa_{A}(t)$ in the thermodynamic limit (i.e., $N\to\infty$ before taking $\gamma\ll 1$). Eq. (21) with $\dot{\kappa}_{A}\sim\gamma$ and $\delta\kappa_{A}^{2}\sim\gamma^{2}t$ implies that the operator-size growth stops at $t\sim\gamma^{-1}$ due to dissipation. The peak of $\kappa_{A}(t)$ is independent of $\gamma$ and $N$. Subsequently, $\kappa_{A}(t)$ gradually approaches the Liouvillian gap $g\sim\gamma$. In this way, there is no plateau regime in the thermodynamic limit, but the amplification of dissipation via the operator spreading occurs, which yields a finite decay rate even for an infinitesimally small $\gamma$.
@@ Acknowledgements.
[…]
${}^{1}$aseda Institute for Advanced Study, Waseda University, Nishi Waseda, Shinjuku-ku, Tokyo 169-0051, Japan
${}^{2}$epartment of Physics, Keio University, Kohoku-ku, Yokohama, Kanagawa 223-8522, Japan
@@ A Bose-Hubbard model
where $\hat{b}_{i}$ and $\hat{b}_{i}^{\dagger}$ are annihilation and creation operators of hard-core bosons at site $i$, respectively. The eigenvalues of the operator $\hat{n}_{i}=\hat{b}_{i}^{\dagger}\hat{b}_{i}$ are 00 and $1$. The periodic boundary condition is assumed. The parameters of the Hamiltonians are given by $(h,h^{\prime},V,V^{\prime})=(1,0.24,1,0.24)$, where the eigenstate thermalization hypothesis has been numerically shown [28]. The Lindblad operators describe a dephasing process and the strength is set to $\gamma=0.01$. The half-filled sector (i.e., $\sum_{i=1}^{N}\hat{n}_{i}=N/2$) is considered, and the steady state has a uniform distribution in this sector.
@@ B Derivation of Equation 21
$=$ $\frac{\braket{\hat{A}(t),[\tilde{\mathcal{L}},\tilde{\mathcal{L}}_{s}]\hat{A}(t)}_{\mathrm{ss}}}{\braket{\hat{A}(t),\hat{A}(t)}_{\mathrm{ss}}}-2\left(\frac{\braket{\hat{A}(t),\tilde{\mathcal{L}}_{s}^{2}\hat{A}(t)}_{\mathrm{ss}}}{\braket{\hat{A}(t),\hat{A}(t)}_{\mathrm{ss}}}-\frac{\braket{\hat{A}(t),\tilde{\mathcal{L}}_{s}\hat{A}(t)}_{\mathrm{ss}}^{2}}{\braket{\hat{A}(t),\hat{A}(t)}_{\mathrm{ss}}^{2}}\right)=\dot{\kappa}_{A}-2\delta\kappa_{A}^{2}.$ (S3)
For the first term of the last expression of Eq. S3, it is plausible to approximate $\dot{\kappa}_{A}$ by
$\dot{\kappa}_{A}^{(0)}=\frac{\braket{\hat{A}^{(0)}(t),[\tilde{\mathcal{L}},\tilde{\mathcal{L}}_{s}]\hat{A}^{(0)}(t)}_{\mathrm{ss}}}{\braket{\hat{A}^{(0)}(t),\hat{A}^{(0)}(t)}_{\mathrm{ss}}}\text{ with }\hat{A}^{(0)}(t)=e^{i\hat{H}t}\hat{A}e^{-i\hat{H}t},$ (S4)
which is obtained by replacing $\hat{A}(t)$ in $\dot{\kappa}_{A}$ by $\hat{A}^{(0)}(t)$. Indeed, our numerical calculation presented in LABEL:fig:derivative (a) shows the validity of this approximation at arbitrary $t$ for small $\gamma$. Thus, $\dot{\kappa}_{A}$ in Eq. S3 expresses the rate of operator spreading due to the unitary dynamics. In contrast, $-2\delta\kappa_{A}^{2}$ corresponds to suppression of the operator growth due to dissipation. The quantity $\delta\kappa_{A}^{2}$ roughly gives the variance of the operator size.
In the acceleration regime, LABEL:fig:derivative (b) and (d) imply $\dot{\kappa}_{A}\propto\gamma$ and $\delta\kappa_{A}^{2}\sim\gamma^{2}t$, respectively. LABEL:fig:derivative (c) and (e) imply both $\dot{\kappa}_{A}$ and $\delta\kappa_{A}^{2}$ are independent of the system size in the acceleration regime. This result is consistent with the recent studies on random unitary circuits [23, 24]. The plateau regime corresponds to the region of time in which $\dot{\kappa}_{A}/\gamma\approx 0$, where $\delta\kappa_{A}^{2}$ is also saturated (see LABEL:fig:derivative (e)). LABEL:fig:derivative (f) and (g) show that $\delta\kappa_{A}^{2}\sim\gamma^{2}N$ in this regime. Finally, the system reaches the asymptotic regime, where $\dot{\kappa}_{A}=2\delta\kappa_{A}^{2}$ and $\kappa_{A}(t)$ remains constant.

### v3

@@ Abstract
which is expressed as ~$\tilde{\mathcal{L}}^{*}\hat{A}=\mathcal{L}(\hat{A}\rho_{\mathrm{ss}})\rho_{\mathrm{ss}}^{-1}$ [17]. We assume that $\rho_{\mathrm{ss}}$ is invertible for simplicity. The symmetrized Liouvillian is a non-positive and Hermitian superoperator (i.e., ~$\tilde{\mathcal{L}}_{s}\leq 0$ and ~s=~$\tilde{\mathcal{L}}_{s}^{*}=\tilde{\mathcal{L}}_{s}$). The non-positivity is shown by using the complete positivity of $\mathcal{L}$. We sort the eigenvalues of ~$\tilde{\mathcal{L}}_{s}$ denoted by $\{s_{\alpha}\}$ in the descending order: $0=s_{0}\geq s_{1}\geq s_{2}\geq\dots$. ~$\tilde{\mathcal{L}}_{s}$ has a zero eigenvalue: ~$\tilde{\mathcal{L}}_{s}\hat{I}=0$, where $\hat{I}$ is the identity operator. The spectral gap $g_{s}$ of ~$\tilde{\mathcal{L}}_{s}$ is defined by
@@
We assume weak dissipation and set $\gamma=0.01$. This open quantum system has been implemented on Rydberg atoms under laser driving and dissipation [20, 21]. Each Rydberg atom is regarded as a two-level system, where the spin up (down) corresponds to the Rydberg (ground) state. In this model, the Liouvillian gap is independent of the system size (i.e., $g\sim O(N^{0})$. See Fig. 1.).
@@
$\kappa_{A}^{(0)}(t)=-\frac{\braket{\hat{A}_{0}(t),\tilde{\mathcal{L}}_{s}\hat{A}_{0}(t)}_{\mathrm{ss}}}{\braket{\hat{A}_{0}(t),\hat{A}_{0}(t)}_{\mathrm{ss}}},\quad\hat{A}_{0}(t)=e^{i\hat{H}t}\hat{A}e^{-i\hat{H}t}.$ (20)
@@
$\frac{d\kappa_{A}(t)}{dt}=v_{A}-2\delta\kappa_{A}^{2},$ (21)
@@
$v_{A}$ $=\frac{\braket{\hat{A}(t),[\tilde{\mathcal{L}},\tilde{\mathcal{L}}_{s}]\hat{A}(t)}_{\mathrm{ss}}}{\braket{\hat{A}(t),\hat{A}(t)}_{\mathrm{ss}}},$
@@
This formula generalizes the result in [25] to any open quantum systems obeying Lindblad equations. $v_{A}/\gamma$ represents how fast the operator size of $A(t)$ spreads with time and $(\delta\kappa_{A}/\gamma)^{2}$ roughly gives the variance of the operator size. Note that the time evolution of $A(t)$ in $v_{A}$ is well approximated by the unitary dynamics without dissipation [18].
The formula accurately explains the observed three dynamic regimes in Fig. 1. In the acceleration regime, recent studies on random unitary circuits [23, 24] indicate $v_{A}\sim\gamma$ and $\delta\kappa_{A}^{2}\sim\gamma^{2}t$. Since $v_{A}\gg\delta\kappa_{A}^{2}$ for small $\gamma$,
@@
The system enters the plateau regime when the operator spreads across the entire system at $t\sim N$. Accordingly, $\kappa_{A}(t)$ is proportional to $\gamma N$. In this regime, $v_{A}\approx 0$, whereas $\delta\kappa_{A}^{2}\sim\gamma^{2}N$ [18]. Thus, during the duration time $\Delta t$ in the plateau regime, $\kappa_{A}(t)$ reduces by the amount proportional to $\gamma^{2}N\Delta t$. This decrease is negligible compared with the plateau value $\sim\gamma N$ up to $\Delta t\sim\gamma^{-1}$. We therefore conclude that $\kappa_{A}(t)$ approximately remains constant up to $t\sim\gamma^{-1}$:
@@
Here, we briefly explain the dynamics of $\kappa_{A}(t)$ in the thermodynamic limit (i.e., $N\to\infty$ before taking $\gamma\ll 1$). Equation 21 with $v_{A}\sim\gamma$ and $\delta\kappa_{A}^{2}\sim\gamma^{2}t$ implies that the operator-size growth stops at $t\sim\gamma^{-1}$ due to dissipation. The peak of $\kappa_{A}(t)$ is independent of $\gamma$ and $N$. Subsequently, $\kappa_{A}(t)$ gradually approaches the Liouvillian gap $g\sim\gamma$. In this way, there is no plateau regime in the thermodynamic limit, but the amplification of dissipation via the operator spreading occurs, which yields a finite decay rate even for an infinitesimally small $\gamma$.
@@ Acknowledgements.
Tatsuhiko Shirai1 and Takashi Mori2
1Waseda Institute for Advanced Study, Waseda University, Nishi Waseda, Shinjuku-ku, Tokyo 169-0051, Japan
2Department of Physics, Keio University, Kohoku-ku, Yokohama, Kanagawa 223-8522, Japan
@@ A Bose-Hubbard model
where $\hat{b}_{i}$ and $\hat{b}_{i}^{\dagger}$ are annihilation and creation operators of hard-core bosons at site $i$, respectively. The eigenvalues of the operator $\hat{n}_{i}=\hat{b}_{i}^{\dagger}\hat{b}_{i}$ are 00 and $1$. The periodic boundary condition is assumed. The parameters of the Hamiltonian are given by $(h,h^{\prime},V,V^{\prime})=(1,0.24,1,0.24)$, where the eigenstate thermalization hypothesis has been numerically shown [28]. The Lindblad operators describe a dephasing process and the strength is set to $\gamma=0.01$. The half-filled sector (i.e., $\sum_{i=1}^{N}\hat{n}_{i}=N/2$) is considered, and the steady state has a uniform distribution in this sector.
@@ B Derivation of Equation 21
$=$ $\frac{\braket{\hat{A}(t),[\tilde{\mathcal{L}},\tilde{\mathcal{L}}_{s}]\hat{A}(t)}_{\mathrm{ss}}}{\braket{\hat{A}(t),\hat{A}(t)}_{\mathrm{ss}}}-2\left(\frac{\braket{\hat{A}(t),\tilde{\mathcal{L}}_{s}^{2}\hat{A}(t)}_{\mathrm{ss}}}{\braket{\hat{A}(t),\hat{A}(t)}_{\mathrm{ss}}}-\frac{\braket{\hat{A}(t),\tilde{\mathcal{L}}_{s}\hat{A}(t)}_{\mathrm{ss}}^{2}}{\braket{\hat{A}(t),\hat{A}(t)}_{\mathrm{ss}}^{2}}\right)=v_{A}-2\delta\kappa_{A}^{2}.$ (S3)
For the first term of the last expression of Eq. S3, it is plausible to approximate $v_{A}$ by
$v_{A}^{(0)}=\frac{\braket{\hat{A}^{(0)}(t),[\tilde{\mathcal{L}},\tilde{\mathcal{L}}_{s}]\hat{A}^{(0)}(t)}_{\mathrm{ss}}}{\braket{\hat{A}^{(0)}(t),\hat{A}^{(0)}(t)}_{\mathrm{ss}}}\text{ with }\hat{A}^{(0)}(t)=e^{i\hat{H}t}\hat{A}e^{-i\hat{H}t},$ (S4)
which is obtained by replacing $\hat{A}(t)$ in $v_{A}$ by $\hat{A}^{(0)}(t)$. Indeed, our numerical calculation presented in LABEL:fig:derivative (a) shows the validity of this approximation at arbitrary $t$ for small $\gamma$. Thus, $v_{A}$ in Eq. S3 expresses the rate of operator spreading due to the unitary dynamics. In contrast, $-2\delta\kappa_{A}^{2}$ corresponds to suppression of the operator growth due to dissipation. The quantity $\delta\kappa_{A}^{2}$ roughly gives the variance of the operator size.
In the acceleration regime, LABEL:fig:derivative (b) and (d) imply $v_{A}\propto\gamma$ and $\delta\kappa_{A}^{2}\sim\gamma^{2}t$, respectively. LABEL:fig:derivative (c) and (e) imply both $v_{A}$ and $\delta\kappa_{A}^{2}$ are independent of the system size in the acceleration regime. This result is consistent with the recent studies on random unitary circuits [23, 24]. The plateau regime corresponds to the region of time in which $v_{A}/\gamma\approx 0$, where $\delta\kappa_{A}^{2}$ is also saturated (see LABEL:fig:derivative (e)). LABEL:fig:derivative (f) and (g) show that $\delta\kappa_{A}^{2}\sim\gamma^{2}N$ in this regime. Finally, the system reaches the asymptotic regime, where $v_{A}=2\delta\kappa_{A}^{2}$ and $\kappa_{A}(t)$ remains constant.